# Degradation-Aware All-in-One Image Restoration via Latent Prior Encoding

### v2

@@ 3.2 Degradation-Aware Reconstruction
Notably, our proposed 3WD achieves linear computational complexity of $O(HW)$ per stage, in contrast to the quadratic complexity $O(H^{2}W^{2})$ of standard attention mechanisms Vaswani et al. (2017). This efficiency stems from utilizing element-wise multiplication ($\odot$) between tensors of shape $H\times W\times C$, requiring exactly $HW\times C$ operations. We perceived final reconstruction with learned features and global residual: $\hat{\mathbf{I}}=\tanh(W_{\text{rec}}D_{1})+\mathbf{I}_{\text{deg}}$.
@@ 4.1 Setup
We evaluated our method under two degradation scenarios: common (non-overlapping) degradations and compound (overlapping) settings. For common degradations, we tested on standard tasks/datasets, including dehazing (SOTS Li et al. (2018)), deraining (Rain100L Fu et al. (2017)), desnowing (CCD Li et al. (2020)), deblurring Nah et al. (2017), deblurring (GoPro Nah et al. (2017)), denoising (DIV2K Agustsson & Timofte (2017)) with random noise levels ($\sigma\in[0,50]$) to improve generalization, and LLIE using the LSD dataset Sharif et al. (2025), captured in uncontrolled environments. In compound degradation, we employed the CDD dataset Guo et al. (2024), combining five degradations (e.g., haze+rain, low-light+haze+snow). Subsets like haze+snow and low-light+rain were reserved for testing generalization comparison on unseen degradations. Ablation and analysis, we combined four common tasks (denoising, desnowing, deraining, and LLIE) to analyze consistency.
We compared DAIR against transformer-based baseline methods (Uformer Wang et al. (2022), Restormer Zamir et al. (2022)), SOTA AIR approaches (ADAIR Cui et al. (2024), AIRNet Li et al. (2022)), prompting-based methods (DFPIR Tian et al. (2025a), PromptIR Vaishnav et al. (2023)), and diffusion-based latent enhancement (DiffUIR Zheng et al. (2024)). Single-task benchmarks included deraining (HDCWNet Zhu et al. (2021), TransWeather Valanarasu et al. (2022)), Retinex-based LLIE (RetinexNet Wei et al. (2018), Diff-Retinex Yi et al. (2023)).
@@
We first pre-train our latent encoder for 200,000 steps, combining six single and five compound (known) degradations using a composite loss (self-reconstruction + KL Kingma & Welling (2014) + discriminative latent regularizer Guo et al. (2023)). The resulting encoder is frozen and reused for all single, compound, and unseen degradation experiments without any fine-tuning. We only tune the main network for TS settings using the combination of L1 and SSIM losses for fair comparison. This network has been trained for 100,000 to 500,000 steps, depending on task complexity. We trained both the latent encoder and the main restoration network using the Adam optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, and learning rate $1\text{e-}4$. All training was conducted on a single NVIDIA RTX 3060 GPU with a batch size of 4, using randomly cropped $256\times 256$ patches as input. All baseline methods are trained using their official implementations and suggested hyperparameters to ensure fairness.
@@ 4.2.1 Multi-task (6D) Restoration
Table 1: Performance comparison across different image restoration tasks. Best results in bold red, second best underlined, and increment over best performing method highlighted in blue.
@@
PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
Uformer Wang et al. (2022) 20.63 43.86 12.10/0.4897 29.15/0.9727 27.55/0.8201 24.80/0.8832 23.67/0.8074 28.74/0.8823 24.34/0.8092
Restormer Zamir et al. (2022) 26.13 141.75 15.33/0.6180 28.51/0.9738 20.41/0.5357 22.98/0.8585 25.41/0.8286 25.99/0.8164 23.10/0.7718
AIRNet Li et al. (2022) 8.93 301.27 8.87/0.3655 25.92/0.9663 28.87/0.8789 19.08/0.7300 22.81/0.8067 24.92/0.8027 21.75/0.7583
PromptIR Vaishnav et al. (2023) 35.59 158.14 14.73/0.5374 30.41/0.9580 23.96/0.6378 27.08/0.9169 26.04/0.8366 30.01/0.8992 25.37/0.7977
DiffUIR Zheng et al. (2024) 36.26 4400.00 10.76/0.4101 30.05/0.9797 20.25/0.6859 20.39/0.8014 25.03/0.8474 27.46/0.8917 22.32/0.7694
ADAIR Cui et al. (2024) 28.78 147.18 12.76/0.4974 30.15/0.9783 28.65/0.8715 25.86/0.8812 25.14/0.8200 29.56/0.8901 25.35/0.8231
DFPIR Tian et al. (2025a) 31.07 151.07 14.84/0.5232 28.29/0.9709 27.78/0.8341 21.74/0.7935 26.10/0.8271 28.02/0.8372 24.46/0.7977
DAIR (Ours) 18.08 45.65 16.87/0.6661 34.08/0.9864 29.12/0.8971 31.50/0.9570 26.82/0.8793 30.51/0.9117 28.15/0.8829
Improvement -10.70 -96.10 +1.54/+0.0481 +3.67/+0.0067 +0.25/+0.0182 +4.42/+0.0401 +0.72/+0.0427 +0.50/+0.0125 +2.78/+0.0598
Figure 3: Visual comparison for six common degradation settings. The proposed method produces consistent and plausible images compared to the existing methods.
Table 1 and Figure 3 illustrate DAIR’s performance across six common degradations. Achieving the highest average PSNR (28.15) and SSIM (0.8829), DAIR delivers significant improvements, including +5.64 PSNR and +0.0758 SSIM in desnowing, and +1.54 PSNR and +0.0481 SSIM in lowlight enhancement. Despite requiring only 45G FLOPs, DAIR outperforms computationally intensive methods like ADAIR Cui et al. (2024)(147G FLOPs) while operating without MP, ensuring scalability and adaptability. By leveraging degradation-aware latent priors, DAIR generates robust, high-quality restorations across diverse degradation types, making it ideal for real-world applications like autonomous systems and healthcare imaging.
@@ 4.2.2 Single-task Evaluation: LLIE and Desnowing
Table 2: Performance comparison on desnowing and LLIE tasks. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
@@
Method PSNR↑/SSIM↑ Method PSNR↑/SSIM↑
HDCWNet Zhu et al. (2021) 29.25/0.9171 RetinexNet Wei et al. (2018) 14.19/0.5183
TransWeather Valanarasu et al. (2022) 23.30/0.7631 Diff-Retinex Yi et al. (2023) 15.38/0.5038
Uformer Wang et al. (2022) 28.86/0.9324 Uformer Wang et al. (2022) 13.05/0.5716
Restormer Zamir et al. (2022) 23.21/0.8686 Restormer Zamir et al. (2022) 15.78/0.6184
ADAIR Cui et al. (2024) 28.31/0.9300 ADAIR Cui et al. (2024) 15.19/0.6094
DAIR (Ours) 32.75/0.9632 DAIR (Ours) 16.30/0.6634
Improvement +3.50/+0.0308 Improvement +0.52/+0.0450
We evaluated DAIR on single restoration tasks to demonstrate its adaptability. Table 2 shows DAIR achieves 32.75 PSNR / 0.9632 SSIM for desnowing, outperforming HDCWNet Zhu et al. (2021) by +3.50 PSNR / +0.0308 SSIM. For LLIE, DAIR achieves 16.30 PSNR / 0.6634 SSIM, improving upon Restormer Zamir et al. (2022) by +0.52 PSNR / +0.0450 SSIM. By leveraging degradation-aware latent priors, DAIR can also adapt to varying degradation types without external MP.
@@ 4.3 Results on Compound Degradation
Table 3: Performance comparison across five compound degradation. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
@@
Uformer Wang et al. (2022) 22.49/0.8889 21.37/0.8252 20.61/0.7838 18.64/0.6802 20.31/0.7150 20.68/0.7786
Restormer Zamir et al. (2022) 23.57/0.8825 21.44/0.7753 20.53/0.7265 18.46/0.6494 19.97/0.6664 20.79/0.7400
AIRNet Li et al. (2022) 18.58/0.7977 11.73/0.6905 13.58/0.6320 12.87/0.5898 13.76/0.5759 14.10/0.6572
PromptIR Vaishnav et al. (2023) 21.78/0.8818 20.63/0.8231 19.83/0.7705 18.42/0.6801 20.74/0.7070 20.28/0.7725
DiffUIR Zheng et al. (2024) 18.42/0.8022 14.28/0.7150 15.79/0.6568 15.33/0.6238 12.60/0.6292 15.28/0.6854
ADAIR Cui et al. (2024) 23.51/0.8966 22.78/0.8420 21.06/0.7908 20.46/0.7229 21.35/0.7215 21.83/0.7948
DFPIR Tian et al. (2025a) 20.81/0.8308 19.01/0.7784 18.82/0.7139 15.58/0.6307 19.70/0.6402 18.78/0.7188
DAIR (Ours) 25.25/0.9259 23.15/0.8541 22.03/0.8200 20.81/0.7613 21.79/0.7772 22.61/0.8277
Improvement +1.68/+0.0293 +0.37/+0.0121 +0.97/+0.0292 +0.35/+0.0384 +0.44/+0.0557 +0.78/+0.0329
@@
We evaluated DAIR on compound degradation scenarios. As shown in Table 3, DAIR achieves the highest average PSNR of 22.61 and SSIM of 0.8277, outperforming the next-best method (ADAIR Cui et al. (2024)) by +0.78 PSNR / +0.0329 SSIM. Notable improvements include ”Haze + Rain” (+1.68 PSNR / +0.0293 SSIM) and ”Low + Snow” (+0.44 PSNR / +0.0557 SSIM). DAIR effectively restores overlapping degradations, demonstrating robust performance across complex scenarios. Fig. 4 highlights DAIR’s superior visual quality, producing artifact-free, natural images.
@@ 4.4 Latent prior Guided Unseen Task Restoration
Table 4: Performance comparison on unknown compound degradations. Best results in bold red, second best underlined. Improvement over the best-performing method highlighted in blue.
@@
ADAIR Cui et al. (2024) 16.07/0.8130 21.67/0.7877 18.87/0.8003
PromptIR Vaishnav et al. (2023) 17.84/0.8032 21.15/0.7752 19.50/0.7892
DFPIR Tian et al. (2025a) 14.50/0.7577 21.52/0.7021 18.01/0.7299
@@
The primary motivation for leveraging latent priors is to address unknown degradations while maintaining a clear distinction from known degradations. Fig. 5 highlights the restoration of unknown compound degradations, comparing DAIR with SOTA methods. Our approach effectively handles these degradations, as demonstrated by t-SNE visualization in Fig. 5 (b). Table 4 summarizes the results, showing that our method achieves notable improvements over existing approaches, with average gains of +1.47 dB PSNR and +0.0412 SSIM for unseen degradations.
@@
4.5 Degradation Map
Figure 6: Visualization and restoration with the proposed DM.
Our proposed DM effectively learns spatially correlated regions, enabling the decoder to focus on ”where” to reconstruct the degraded image. As shown in Fig. 6, the comparison with traditional skip connections highlights that our learnable approach precisely identifies affected areas and significantly outperforms conventional methods.
4.6 Restoration with 3WD
We evaluate DAIR with 3WD against self-attention (SA) Vaswani et al. (2017) and cross-attention (CA) Vaswani et al. (2017) baselines. As shown in Table 5, DAIR with 3WD consistently outperforms both baselines, achieving average improvements of +7.72 dB PSNR and +0.2605 SSIM, with significant gains in desnowing (+11.89 dB) and denoising (+8.24 dB). Additionally, DAIR operates with linear complexity $O(d^{2}\times H\times W)$ compared to the baselines’ quadratic $O(d\times H^{2}\times W^{2})$, delivering 85-257× speedup and 76-89% memory reduction while answering our reasoning cues.
Table 5: Performance comparison of proposed 3WD restoration with existing attention mechanism. Best results in bold red, second best underlined. Improvement over the best performing method highlighted in blue.
Method Complexity Denoising Desnowing Deraining Lowlight Average
(Speedup) PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑ PSNR↑/SSIM↑
DAIR + SA Vaswani et al. (2017) $O(H^{2}W^{2})$ 18.90/0.6213 19.70/0.6738 22.94/0.6541 15.16/0.4348 19.18/0.5960
DAIR + CA Vaswani et al. (2017) $O(H^{2}W^{2})$ 20.87/0.6290 16.07/0.6469 20.90/0.6448 13.00/0.3854 17.71/0.5765
DAIR + 3WD (Ours) $O(HW)$ 29.11/0.8971 31.59/0.9581 30.21/0.9070 16.68/0.6638 26.90/0.8565
Improvement Linear scaling +8.24/+0.2681 +11.89/+0.2843 +7.27/+0.2529 +1.52/+0.2290 +7.72/+0.2605

### v3

@@ 3.2 Degradation-Aware Reconstruction
Notably, our proposed 3WD achieves linear computational complexity of $O(HW)$ per stage, in contrast to the quadratic complexity $O(H^{2}W^{2})$ of standard attention mechanisms (vaswani2017attention). This efficiency stems from utilizing element-wise multiplication ($\odot$) between tensors of shape $H\times W\times C$, requiring exactly $HW\times C$ operations. We perceived final reconstruction with learned features and global residual: $\hat{\mathbf{I}}=\tanh(W_{\text{rec}}D_{1})+\mathbf{I}_{\text{deg}}$.
@@ 4.1 Setup
We evaluated our method under two degradation scenarios: common (non-overlapping) degradations and compound (overlapping) settings. For common degradations, we combined six widely used tasks, including dehazing (SOTS (li2018benchmarking)), deraining with heavy rain (Rain100H (fu2017deep)), desnowing (CCD (li2020desnownet)), real-world deblurring (nah2017deep), deblurring (GoPro (nah2017deep)), denoising (DIV2K (agustsson2017div2k)) with random noise levels ($\sigma\in[0,50]$) to improve generalization, and real-world LLIE using the LSD dataset (sharif2025illuminatingdarknesslearningenhance). In compound degradation, we employed the CDD dataset (guo2024onerestore), combining five degradations (e.g., haze+rain, low-light+haze+snow). Subsets like haze+snow and low-light+rain were reserved for testing generalization comparison on unseen degradations. We also included several real-world out-of-distribution datasets, such as SIDD for noisy images, LSD-U for unseen low-light scenarios, underwater image enhancement datasets, and medical image enhancement and denoising datasets.
We compared DAIR against transformer-based baseline methods (Uformer (wang2022uformer), Restormer (zamir2022restormer)), SOTA AIR approaches (ADAIR (cui2024adair), AIRNet (li2022all)), prompting-based methods (DFPIR (tian2025degradation), PromptIR (vaishnav2023promptir)), and diffusion-based latent enhancement (DiffUIR (zheng2024selective)). Single-task benchmarks included deraining (HDCWNet (zhu2021hdcwnet), TransWeather (valanarasu2022transweather)), Retinex-based LLIE (RetinexNet (wei2018deep), Diff-Retinex (yi2023diff)).
@@
We first pre-train our latent encoder for 200,000 steps, combining six single and five compound (known) degradations using a composite loss (self-reconstruction + KL (kingma2014auto) + discriminative latent regularizer (guo2023letting)). The resulting encoder is frozen and reused for all single, compound, and unseen degradation experiments without any fine-tuning. We only tune the main network for TS settings using the combination of L1 and SSIM losses for fair comparison. This network has been trained for 100,000 to 500,000 steps, depending on task complexity. We trained both the latent encoder and the main restoration network using the Adam optimizer with hyperparameters $\beta_{1}=0.9$, $\beta_{2}=0.99$, and learning rate $1\text{e-}4$. All training was conducted on a single NVIDIA RTX 3060 GPU with a batch size of 4, using randomly cropped $256\times 256$ patches as input. All baseline methods are trained using their official implementations and suggested hyperparameters to ensure fairness.
@@ 4.2.1 Multi-task (6D) Restoration
Table 1: Performance comparison across six different image restoration tasks under all-in-one setting: a unified model is trained on a combined set of images obtained from all degradation types and levels. Best results in bold red, second best underlined, and increment over best performing method highlighted in blue.
@@
PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓
Uformer 20.63 43.86 12.10/0.4897/0.3621 29.15/0.9727/0.0195 27.55/0.8201/0.1334 24.80/0.8832/0.0904 23.67/0.8074/0.2401 28.74/0.8823/0.1051 24.34/0.8092/0.1584
Restormer 26.13 141.75 15.33/0.6180/0.2697 28.51/0.9738/0.0154 20.41/0.5357/0.4148 22.98/0.8585/0.1213 25.41/0.8286/0.2080 25.99/0.8164/0.1547 23.10/0.7718/0.1973
AIRNet 8.93 301.27 8.87/0.3655/0.3679 25.92/0.9663/0.0237 28.87/0.8789/0.1020 19.08/0.7300/0.2596 22.81/0.8067/0.2284 24.92/0.8027/0.1692 21.75/0.7583/0.1918
PromptIR 35.59 158.14 14.73/0.5374/0.3665 30.41/0.9580/0.0221 23.96/0.6378/0.2339 27.08/0.9169/0.0634 26.04/0.8366/0.2069 30.01/0.8992/0.0669 25.37/0.7977/0.1600
DiffUIR 36.26 4400.00 10.76/0.4101/0.4290 30.05/0.9797/0.0151 20.25/0.6859/0.2329 20.39/0.8014/0.1709 25.03/0.8474/0.2019 27.46/0.8917/0.0719 22.32/0.7694/0.1870
ADAIR 28.78 147.18 12.76/0.4974/0.3605 30.15/0.9783/0.0135 28.65/0.8715/0.1043 25.86/0.8812/0.0966 25.14/0.8200/0.2307 29.56/0.8901/0.0889 25.35/0.8231/0.1491
DFPIR 31.07 151.07 14.84/0.5232/0.3766 28.29/0.9709/0.0214 27.78/0.8341/0.1433 21.74/0.7935/0.2189 26.10/0.8271/0.2230 28.02/0.8372/0.1506 24.46/0.7977/0.1890
DAIR (Ours) 18.08 45.65 16.87/0.6661/0.2559 34.08/0.9864/0.0100 29.12/0.8971/0.0826 31.50/0.9570/0.0264 26.82/0.8793/0.1653 30.51/0.9117/0.0663 28.15/0.8829/0.1011
Improvement -10.70 -96.10 +1.54/+0.0481/-0.0138 +3.67/+0.0067/-0.0035 +0.25/+0.0182/-0.0194 +4.42/+0.0401/-0.0370 +0.72/+0.0427/-0.0366 +0.50/+0.0125/-0.0006 +2.78/+0.0598/-0.0480
Figure 3: Visual comparison for six common degradation settings under the all-in-one setting. The proposed method produces consistent and plausible images compared to the existing methods.
Table 1 and Figure 3 illustrate DAIR’s performance across six common degradations under all-in-one setting. Achieving the highest average PSNR (28.15) and SSIM (0.8829), DAIR delivers significant improvements, including +5.64 PSNR and +0.0758 SSIM in desnowing, and +1.54 PSNR and +0.0481 SSIM in lowlight enhancement. Despite requiring only 45G FLOPs, DAIR outperforms computationally intensive methods like ADAIR (cui2024adair)(147G FLOPs) while operating without MP, ensuring scalability and adaptability. By leveraging degradation-aware latent priors, DAIR generates robust, high-quality restorations across diverse degradation types, making it ideal for real-world applications like autonomous systems and healthcare imaging.
@@ 4.2.2 Single-task Evaluation: LLIE and Desnowing
Table 2: Performance comparison on desnowing and LLIE tasks.
@@
Method PSNR↑/SSIM↑/LPIPS↓ Method PSNR↑/SSIM↑/LPIPS↓
HDCWNet 29.25/0.9171/0.0517 RetinexNet 14.19/0.5183/0.3812
TransWeather 23.30/0.7631/0.1739 Diff-Retinex 15.38/0.5038/0.3841
Uformer 28.86/0.9324/0.0904 Uformer 13.05/0.5716/0.3621
Restormer 23.21/0.8686/0.1059 Restormer 15.78/0.6184/0.3372
ADAIR 28.31/0.9300/0.0457 ADAIR 15.19/0.6094/0.3325
DAIR (Ours) 32.75/0.9632/0.0213 DAIR (Ours) 16.30/0.6634/0.3279
Improvement +3.50/+0.0308/-0.0244 Improvement +0.52/+0.0450/-0.0046
We evaluated DAIR on single restoration tasks to demonstrate its adaptability. Table 2 shows DAIR achieves 32.75 PSNR / 0.9632 SSIM for desnowing, outperforming HDCWNet (zhu2021hdcwnet) by +3.50 PSNR / +0.0308 SSIM. For LLIE, DAIR achieves 16.30 PSNR / 0.6634 SSIM, improving upon Restormer (zamir2022restormer) by +0.52 PSNR / +0.0450 SSIM. By leveraging degradation-aware latent priors, DAIR can also adapt to varying degradation types without external MP.
@@ 4.3 Results on Compound Degradation
Table 3: Quantitative evaluation of restoration performance on five compound degradation types, derived from combinations of low-light, rain, haze, and snow datasets.
@@
PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓ PSNR↑/SSIM↑/LPIPS↓
Uformer 22.49/0.8889/0.0867 21.37/0.8252/0.1695 20.61/0.7838/0.1908 18.64/0.6802/0.3052 20.31/0.7150/0.2513 20.68/0.7786/0.2007
Restormer 23.57/0.8825/0.0959 21.44/0.7753/0.1884 20.53/0.7265/0.2367 18.46/0.6494/0.3309 19.97/0.6664/0.3003 20.79/0.7400/0.2304
AIRNet 18.58/0.7977/0.1987 11.73/0.6905/0.2587 13.58/0.6320/0.3731 12.87/0.5898/0.4141 13.76/0.5759/0.4372 14.10/0.6572/0.3364
PromptIR 21.78/0.8818/0.1012 20.63/0.8231/0.1850 19.83/0.7705/0.2223 18.42/0.6801/0.3171 20.74/0.7070/0.2645 20.28/0.7725/0.2180
DiffUIR 18.42/0.8022/0.1758 14.28/0.7150/0.2562 15.79/0.6568/0.3308 15.33/0.6238/0.3500 12.60/0.6292/0.2988 15.28/0.6854/0.2823
ADAIR 23.51/0.8966/0.0791 22.78/0.8420/0.1419 21.06/0.7908/0.1774 20.46/0.7229/0.2416 21.35/0.7215/0.2444 21.83/0.7948/0.1769
DFPIR 20.81/0.8308/0.1548 19.01/0.7784/0.2339 18.82/0.7139/0.2800 15.58/0.6307/0.3366 19.70/0.6402/0.3227 18.78/0.7188/0.2656
DAIR (Ours) 25.25/0.9259/0.0637 23.15/0.8541/0.1408 22.03/0.8200/0.1618 20.81/0.7613/0.2233 21.79/0.7772/0.1874 22.61/0.8277/0.1554
Improvement +1.68/+0.0293/-0.0154 +0.37/+0.0121/-0.0011 +0.97/+0.0292/-0.0156 +0.35/+0.0384/-0.0183 +0.44/+0.0557/-0.0570 +0.78/+0.0329/-0.0215
@@
We evaluated DAIR on five compound degradation scenarios derived from combinations of low-light, rain, haze, and snow datasets. As shown in Table 3, DAIR achieves the highest average PSNR of 22.61 and SSIM of 0.8277, outperforming the next-best method (ADAIR (cui2024adair)) by +0.78 PSNR / +0.0329 SSIM. Notable improvements include ”Haze + Rain” (+1.68 PSNR / +0.0293 SSIM) and ”Low + Snow” (+0.44 PSNR / +0.0557 SSIM). DAIR effectively restores overlapping degradations, demonstrating robust performance across complex scenarios. Fig. 4 highlights DAIR’s superior visual quality, producing artifact-free, natural images.
@@ 4.4 Latent prior Guided Unseen Task Restoration
Our latent prior descriptor enables robust handling of unknown degradations while maintaining clear separation from known ones. As shown in Table 5, the VAE-based encoder learns degradation characteristics directly from corrupted images, achieving strong clustering for both seen and unseen degradations with KNN accuracy of 0.994/0.976, separation ratios of 2.911/1.523, and optimal global metrics (lowest Davies-Bouldin index, highest Calinski-Harabasz score). This learned prior guides the main network with degradation-specific descriptions, enabling clear latent separation for different degradation, where existing methods fail (Fig. 5(b)). Consequently, DAIR restores unknown compound degradations effectively (Fig. 5(a)), achieving average gains of +1.47 dB PSNR and +0.0412 SSIM over SOTA methods (Table 5). Please refer appendix for more details.
Table 4: Latent prior descriptor yields strong clustering and separation for seen/unseen degradations.
Model Seen Unseen Overall Global Metrics
Acc↑ Sep↑ Acc↑ Sep↑ Acc↑ Sep↑ DB↓ CH↑
w/o SupCon 0.864 2.00 0.787 1.32 0.874 1.79 2.69 254.48
Steps = 0 0.826 3.11 0.756 1.33 0.844 2.28 4.84 458.34
Steps = 100k 0.871 2.36 0.828 1.27 0.892 1.95 2.45 289.84
Full Training 0.994 2.91 0.976 1.52 0.991 3.22 1.23 549.34
Table 5: Performance comparison on unknown compound degradations.
@@
ADAIR 16.07/0.8130 21.67/0.7877 18.87/0.8003
PromptIR 17.84/0.8032 21.15/0.7752 19.50/0.7892
DFPIR 14.50/0.7577 21.52/0.7021 18.01/0.7299
@@
4.5 Analysis of Reasoning Restoration
4.5.1 Latent Prior Encoding
The distance matrix in Fig. 7 (a) shows aggregated latent representations: strong intra-class consistency (diagonal: 160–194) and clear inter-class separation. Related degradations have moderate distances (e.g., Haze–Lowlight: 102, Haze–Blur: 86), while dissimilar pairs show larger separations (e.g., Snow–Lowlight: 201, Snow–Haze: 136). As shown in Fig. 7 (b), encoder stages learn progressive feature selection based on these latent priors: early stages (L1–L2) show higher distances ( 180–200) indicating stronger luminance preference for structural information, while deeper stages (L3–L4) show lower distances ( 60–100) reflecting increased chrominance preference for color-specific restoration. This degradation-aware hierarchical guidance enables DAIR to adaptively leverage structural and chromatic information across different encoder depths for different degradations, validating our ”which” reasoning.
Figure 6: Latent prior encoding validation. (a) Degradation separability via distance matrix. (b) Encoder progression: structural (L1–L2) to chromatic (L3–L4) features.
Figure 7: Degradation localization. Our learnable map precisely identifies affected regions (yellow box).
4.5.2 Degradation Map
Table 6 quantitatively evaluates our degradation map localization accuracy (”where”). To evaluate degradation localization, we generate ground truth masks by computing pixel-wise absolute differences $|I_{\text{degraded}}-I_{\text{clean}}|$ and compare them with our predicted degradation maps. Our method achieves a mean IoU of 0.861±0.169 and a Dice score of 0.912±0.118. Pixel-level degradations (Noise, Lowlight, Haze) achieve near-perfect localization (IoU >> 0.95), while spatially-varying degradations (Rain, Snow, Blur) maintain strong performance despite higher variance. As shown in Fig. 7, our learnable degradation maps precisely identify affected regions and significantly outperform traditional skip connections, enabling the decoder to focus reconstruction on degraded areas.
Table 6: Quantitative evaluation of degradation map localization (”where”) using IoU and Dice metrics across pixel-wise and spatially-varying degradations.
Metric Noise Lowlight Haze Rain Snow Blur Mean ± Std
[…]
4.5.3 Latent Encoded Fusion
Table 7: Latent Prior Encoding Quality: Quantitative evaluation of degradation-aware representation learning across different degradation types
Metric Blur Haze Lowlight Noise Rain Snow GLOBAL
[…]
Table 7 validates our latent fusion for degradation semantic encoding (”what”). By fusing luminance and chrominance representations with global priors via adaptive modulation, our method achieves a separation ratio of 1.56 and KNN accuracy of 0.932. Frequency-domain degradations (blur, noise) show perfect classification (Acc = 1.0), while pixel-level and spatially-varying degradations maintain strong separability (Acc >> 0.88, Sep >> 1.15). These results demonstrate that our latent fusion captures degradation-specific characteristics without explicit prompts, guiding the decoder in determining ”what” content to reconstruct.
4.5.4 Restoration with 3WD
Table 8: Comparison and restoration with 3WD.
Method Complexity Memory (MB)↓ FPS↑ PSNR↑/SSIM↑
CA $O(H^{2}W^{2})$ - - 17.71/0.5765
SA $O(H^{2}W^{2})$ - - 19.05/0.5444
WA $O(M^{2}\cdot HW)$ 4519.90 8.77 23.42/0.7944
3WD (Ours) $O(HW)$ 2045.94 45.69 26.90/0.8565
Improvement Linear -54.7% +5.2× +3.48/+0.0621
We evaluate DAIR with 3WD for leveraging reasoning cues against attention-based baselines: self-attention (SA), cross-attention (CA) (vaswani2017attention), and window attention (WA) (liang2021swinir). As shown in Table 8, our method achieves 26.90 dB PSNR and 0.8565 SSIM, outperforming the best baseline (WA) by +3.48 dB and +0.0621 SSIM. Critically, our linear complexity $O(HW)$ enables 5.2× speedup and 54.7% memory reduction compared to WA’s quadratic $O(M^{2}\cdot HW)$ complexity. We achieve 45.69 FPS with 4593.76 MB at HD 720p, with similar efficiency maintained at 2K resolution (20.45 FPS, 4573.29 MB). This validates that our 3WD reasoning mechanism effectively captures spatial-degradation dependencies without quadratic computational overhead.